a juggler keeps throwing balls up in the air one after the other at regular intervals,throwing the second ball 
when the first is at its highest point.how high do the ball rise if he throws four balls in one second?

Grade:10

1 Answers

Rishu Garg
6 years ago
Time of ball at half of maximum height is 1sec.Time at maximum height is 2sec.V=u+gt0=u-10(2)U=20m/su^2=2gs20×20=2×10×sMaximum height =20m